OpenStudy (anonymous):

Find the length of a side of rhombus ABCD if AE=6".

OpenStudy (anonymous):



OpenStudy (mathstudent55):

What do you know about the diagonals of a rhombus?

OpenStudy (anonymous):

AD=12"

OpenStudy (anonymous):

CB=2x

OpenStudy (mathstudent55):

Yes, the diagonals of a rhombus bisect each other, but there is another important property of the diagonals of a rhombus.

OpenStudy (anonymous):

What is the other property?

OpenStudy (mathstudent55):

The are perpendicular to each other.

OpenStudy (mathstudent55):

Also, all sides of a rhombus are congruent. 

OpenStudy (mathstudent55):

Now you can use the Pythagorean theorem to find x.

OpenStudy (anonymous):

a^2+b^2=c^2. So, 2x^2-6^2=x^2.

OpenStudy (mathstudent55):

Yes, you have the right idea, but you need to be careful when you write it.

OpenStudy (mathstudent55):

\((2x)^2 - 6^2 = x^2\) The parentheses are important because you will get 4x^2. Without parentheses you have 2x^2 which is incorrect.

OpenStudy (mathstudent55):

I'd start with \(x^2 + 6^2 = (2x)^2\) Then change it to: \((2x)^2 - x^2 = 6^2\) Now continue from there.

OpenStudy (mathstudent55):

Did you get an answer?

OpenStudy (anonymous):

9

OpenStudy (mathstudent55):

???

OpenStudy (mathstudent55):

The next step is \(4x^2 - x^2 = 36\) \(3x^2 = 36\) Divide both sides by 3 and then take the square root.

OpenStudy (anonymous):

oh opps i did something else

OpenStudy (anonymous):

x^2=12. So, x=3.4641016

OpenStudy (mathstudent55):

No don't find a decimal approximation. \(x^2 = 12\) \(x = \sqrt {12}\) \(x = \sqrt {4 \times 3} \) \(x = 2\sqrt 3\) Now we have x. The side of the rhombus is 2x, so multiply x by 2.

OpenStudy (anonymous):

Side= \[4\sqrt{3}\approx6.93\]

OpenStudy (anonymous):

I understand now. Thanks for the help @mathstudent55 :D Much appreciated.

OpenStudy (mathstudent55):

Excellent work! You're welcome.